140 g of S2Cl2 was collected from the reaction of 640 g of Sulfur and 142 g of Chlorine. How much of the excess reactant was left over?


1
Expert's answer
2022-06-09T05:02:05-0400

Solution:

The molar mass of sulfur (S8) is 256 g/mol

The molar mass of chlorine gas (Cl2) is 71 g/mol


Calculate moles of each reactant:

Moles of S8 = (640 g S8) × (1 mol S8 / 256 g S8) = 2.5 mol S8

Moles of Cl2 = (142 g Cl2) × (1 mol Cl2 / 71 g Cl2) = 2.0 mol Cl2


Balanced chemical equation:

S8(l) + 4Cl2(g) → 4S2Cl2(l)

According to stoichiometry:

4 mol of Cl2 react with 1 mol of S8

Thus, 2.0 mol of Cl2 react with:

(2.0 mol Cl2) × (1 mol S8 / 4 mol Cl2) = 0.5 mol S8

However, initially there is 2.5 mol of S8 (according to the task)

Therefore, Cl2 acts as limiting reactant and S8 is excess reactant


Sulfur (S8) is excess reactant

Therefore,

(2.5 mol S8 − 0.5 mol S8) = 2.0 mol S8 – excess


The molar mass of sulfur (S8) is 256 g/mol

Therefore,

(2.0 mol S8) × (256 g S8 / 1 mol S8) = 512 g S8

512 grams of S8 were left over


Answer: 512 grams of the excess reactant (S8) was left over